OpenStudy (anonymous):

this is a fraction question plz show your work 4/5p-2/7p

OpenStudy (id21):

\[\frac{4}{5p}-\frac{2}{7p}=\frac{4 \times 7 -2 \times 5}{5\times 7p}=\frac{18}{35p}\]

OpenStudy (anonymous):

omg thx u helped so much